Circular No. 6656

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 L. Piro, E. Costa, M. Feroci, and P. Soffitta, Istituto di Astrofisica

Spaziale, CNR, Frascati; L. A. Antonelli, F. Fiore, and P. Giommi, BeppoSAX, Rome; A. Owens and A. Parmar, Space Science Department, ES-TEC; S. Molendi, Istituto di Fisica Cosmica e Tecnologie Relative, CNR, Milano; G. Cusumano, Istituto di Fisica Cosmica e Applicazioni Informatica, CNR, Palermo; J. in ’t Zand and J. Heise, Space Research Organization of the Netherlands, Utrecht; F. Frontera and G. Zavattini, Università di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; M. Smith, G. Gandolfi, V. Torroni, G. Spoliti, A. Coletta, M. Capalbi, S. Rebecchi, D. Ricci, L. Bruca, and G. Crisigiovanni, BeppoSAX, Rome; L. Salotti, G. Gennaro, and C. DeLibero, Nuova Telespazio, Rome; and R. C. Butler, Agenzia Spaziale Italiana, Rome, report: “The BeppoSAX GRBM/WFC error box of GRB 970508 was observed with the BeppoSAX Narrow Field Instruments (NFI) on May 9.1375 UT for ∼ 36 000 s (∼ 5.7 hr after the γ-ray burst). A previously unknown x-ray source, 1SAX J0653.8+7916, has been detected by the MECS (units 2 and 3) and LECS at α = 6^{h}53^{m}46^{s}.7, δ = +79^{o}16′02′′ (equinox 2000.0; estimated error radius 50′′). This position lies in the BeppoSAX WFC error box (IAUC 6649, 6654). The previously known source, 1RXS J065213.8+790855, lying outside the WFC error box, is also detected at α = 6^{h}52^{m}14^{s}.8, δ = +79^{o}09′16′′, consistent within the errors with the ROSAT measurement. The countrates of 1SAX J0653.8+7916 were 0.0070 ± 0.0007 and 0.0043 ± 0.0006 count/s in the MECs (2–10 keV) and LECS (0.5–5 keV), respectively, corresponding to fluxes of (6.3 ± 0.6) and (7.0 ± 0.8) × 10^{−13} erg cm^{−2} s^{−1}. The optical transient reported on IAUC 6654 coincides, within the errors, with 1SAX J0653.8+7916. A further observation with the BeppoSAX NFI is foreseen on May 11 to study the temporal evolution.”

A. O. Jaunsen, and T. Grav, Institute for Theoretical Astrophysics, Oslo; R. Oestensen, Nordic Optical Telescope (NOT); H. Pedersen, Copenhagen University Observatory; and Piro, Feroci, Costa, Frontera, Palazzi, and Heise report: “Data taken at the 2.56-m NOT confirms the variability of the source reported by Bond (IAUC 6654). On May 10.92 UT, we found V = 19.65 for the candidate, V = 20.10 for the nearby star; these magnitudes were calibrated against GSC 4534.1515 (V = 13.3). Unfiltered images taken during the two preceding nights further document the rise in brightness.”

1997 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green